# Counting Hallucinations in Diffusion Models

### v1

@@ Abstract
Diffusion probabilistic models (DPMs) have demonstrated remarkable progress in generative tasks, such as image and video synthesis. However, they still often produce hallucinated samples (hallucinations) that conflict with real-world knowledge, such as generating an implausible duplicate cup floating beside another cup. Despite their prevalence, the lack of feasible methodologies for systematically quantifying such hallucinations hinders progress in addressing this challenge and obscures potential pathways for designing next-generation generative models under factual constraints. In this work, we bridge this gap by focusing on a specific form of hallucination, which we term counting hallucination, referring to the generation of an incorrect number of instances or structured objects, such as a hand image with six fingers, despite such patterns being absent from the training data. To this end, we construct a dataset suite CountHalluSet, with well-defined counting criteria, comprising ToyShape, SimObject, and RealHand. Using these datasets, we develop a standardized evaluation protocol for quantifying counting hallucinations, and systematically examine how different sampling conditions in DPMs, including solver type, ODE solver order, sampling steps, and initial noise, affect counting hallucination levels. Furthermore, we analyze their correlation with common evaluation metrics such as Fréchet Inception Distance (FID), revealing this widely used image quality metric fail to capture counting hallucinations consistently. This work aims to take the first step toward systematically quantifying hallucinations in diffusion models and offer new insights into the investigation of hallucination phenomena in image generation.
@@ 1 Introduction
Diffusion probabilistic models (DPMs) (Ho et al., 2020; Dhariwal & Nichol, 2021; Rombach et al., 2022), commonly referred to as diffusion model, have achieved considerable success in generative modeling, producing high-quality samples across various tasks such as text-to-image generation (Ho & Salimans, 2021; Rombach et al., 2022), audio synthesis (Kong et al., 2021), video generation (Ho et al., 2022), 3D object generation (Luo & Hu, 2021; Vahdat et al., 2022), and protein structure prediction (Jumper et al., 2021; Abramson et al., 2024). Despite these impressive advancements, diffusion models still struggle to faithfully capturing factual properties from training data, often generating outputs that violate consistency with real-world knowledge. A notable example is the generation of objects with an incorrect number of spatially intricate components, such as extra human fingers (Narasimhaswamy et al., 2024) or additional animal limbs (Borji, 2023).
Recent studies have increasingly focused on identifying and analyzing common failure modes in text-guided diffusion models. For instance, Borji (2023) systematically concluded several qualitative shortcomings in text-to-image models. Additionally, Liu et al. (2024) proposed an adversarial search method to uncover undesirable behaviors and failure cases in text-to-image generation, while Wu et al. (2023) employed an iterative approach to discover and mitigate spurious correlations.
More recently, hallucination, a specific failure mode in diffusion models, has garnered attention in the research community (Aithal et al., 2024). Unlike previously discussed text‑prompt‑based failure modes, which can be evaluated by measuring the alignment between the input conditions and outputs (e.g., CLIP score (Hessel et al., 2021) and VQA score (Huang et al., 2025; Zarei et al., 2024)), or more obvious failure modes such as severe distortions, blur, and artifacts, hallucinated samples (hallucinations) often appear visually plausible, and emerge independently of the conditioning signal, manifesting as subtle violations of factual consistency with respect to the training data, such as incorrect object counts, geometries that appear plausible but are physically impossible, or other breaches of physical laws. These fine-grained distinctions between valid samples and hallucinations present a critical challenge for current AI systems, which often fail to detect them reliably. For example, even state-of-the-art large vision-language models like GPT-4o still struggle with basic tasks like accurately counting the number of fingers in a clear hand image. To this end, establishing a standardized and well-defined protocol for reliably quantifying hallucinated samples is essential, as it provides a foundation for outlining pathways toward designing next-generation generative models with strict factual consistency, an ability that is critical for their role as world models (Ha & Schmidhuber, 2018; Ding et al., 2024; Bruce et al., 2024).
@@
In our quantification protocol for counting hallucinations, a critical step is the accurate identification and quantification of objects. To this end, we employ counting models tailored to each dataset. Leveraging these datasets and counting models, we examine how various denoising conditions (e.g., sampling steps, initial noise, the order of ODE solvers, and solver type) within diffusion models affect the counting hallucination rate. Besides, we conduct correlation analysis between counting hallucination rate and Fréchet Inception Distance (FID).
@@
  • Commonly applied numerical strategies in diffusion models, such as increasing sampling steps from 25 to 100, can help mitigate counting hallucinations in synthetic datasets but exacerbates them in the RealHand dataset.
  • The correlations between counting hallucination rate and FID are dataset-dependent and solver-dependent, rather than intrinsic.
  • Based on one of our findings, we propose a simple but effective method, termed joint-diffusion models, which can significantly reduce counting-based hallucinations and non-counting failures in the RealHand dataset.
@@ 2.1 Hallucinations in Diffusion Models
Inspired by the growing attention to hallucination in large language models (LLMs) (Kadavath et al., 2022; Ji et al., 2023; Huang et al., 2023), the study of hallucinations in unconditional diffusion models has also garnered increasing interest. In particular, Aithal et al. (2024) proposed a qualitative hypothesis suggesting that hallucinated samples may arise from data interpolating between distinct modes. Despite this insight into potential causes, a systematic quantitative analysis of hallucinations in diffusion models remains lacking.
@@ 2.2 Quantitative Metrics in Image Generation
The Visual Question Answering (VQA) score (Huang et al., 2025; Zarei et al., 2024) and the CLIP score (Hessel et al., 2021) are two widely used metrics for evaluating the performance of text‑guided generated images. However, such evaluation methods that rely on language‑based models are inherently limited when applied to assess hallucinated images. One line of evidence is the statistical lower bound on the propensity of pretrained language models to hallucinate certain fact types, which constrains the upper bound of detection accuracy based solely on language models (Kalai & Vempala, 2024). Another indication is the persistent performance gap between human annotators and LLMs in classifying hallucinated textual content (Lin et al., 2022; Li et al., 2023).
In addition, several widely used quantitative evaluation metrics, such as FID (Heusel et al., 2017), IS (Salimans et al., 2016), and Precision & Recall (Sajjadi et al., 2018), are commonly employed in image generation tasks. These metrics rely on the feature embeddings extracted from a pretrained Inception-v3 model (Szegedy et al., 2016). However, whether these metrics can capture the severity of hallucinations in generated datasets remains an open question. In this work, we aim to address this gap from the perspective of counting hallucinations.
@@ 3.1 Forward (Diffusion) Process
Forward Process. Given the observed data ${\bm{x}}_{0}\sim q_{0}({\bm{x}}_{0})$, Diffusion Probabilistic Models (DPMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020) introduce a forward process (a.k.a, diffusion process) that gradually perturbs the data by adding Gaussian noise over $T$ discrete timesteps. At at each timestep $t$, the marginal distribution of the noisy sample ${\bm{x}}_{t}$ conditioned on the original data ${\bm{x}}_{0}$ follows:
@@
where $t\in\{0,1,\ldots,T\}$ and $\{\beta_{i}\}_{i=1}^{T}$ is a pre-defined noise schedule dependent on $t$, commonly chosen to be linear (Ho et al., 2020) or cosine (Nichol & Dhariwal, 2021) schedules. The factor $\bar{\alpha}_{t}$ controls the signal-to-noise ratio at step $t$: $\bar{\alpha}_{0}=1$ (no noise) and $\bar{\alpha}_{T}\approx 0$ (nearly pure noise). An equivalent formulation of Eq. (1) is:
@@
As $T$ becomes sufficiently large, the marginal distribution $q_{0:T}({\bm{x}}_{T}|{\bm{x}}_{0})$ approaches a standard Gaussian distribution ${\mathcal{N}}(\bm{0},{\bm{I}})$. In widely-used DPMs (Ho et al., 2020; Nichol & Dhariwal, 2021), $T$ is typically set to 1000.
@@
Training Objective. The reverse process (a.k.a., denoising process) of DPMs aims to iteratively recover the observed data based on the diffused data ${\bm{x}}_{T}$. Specifically, DPMs attempt to predict the noise $\bm{\epsilon}$} from the data ${\bm{x}}_{t}$ by using a neural network ${\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)$, known as noise prediction model. The parameter $\theta$ is optimized by minimizing the following objective (Ho et al., 2020; Song et al., 2022):
@@
where $\nabla_{{\bm{x}}_{t}}\,log\,q_{t}({\bm{x}}_{t})$ is the only unknown term, referred as score function (Song et al., 2022).
@@
ODE-Based Accelerated Sampling. The ancestral sampling in Eq. (4) contains a stochastic term (Kloeden et al., 1992), while there is an associated deterministic process, which can be formulated as an ordinary differential equation (ODE) (a.k.a, probability flow ODE (Song et al., 2022)):
@@
where ${\mathcal{G}}(t_{2},t_{1})=e^{\int_{t_{1}}^{t_{2}}f(v)\text{d}v}$ is the state transition operator for the linear part $\frac{\text{d}{\bm{z}}}{\text{d}\tau}=f(\tau){\bm{z}}$. $\Delta t>0$ denotes the step size, allowing ODE-based solvers to accelerate the sampling process with larger step sizes compared to ancestral sampling. The total number of sampling steps is given by $T/\Delta t$. Since the linear part ${\mathcal{G}}(t_{2},t_{1})$ can be computed in closed form, most accelerated ODE-based samplers (Song et al., 2021; Liu et al., 2022; Zhang & Chen, 2023; Lu et al., 2022; 2023) devote their numerical efforts, such as higher solver orders and more sampling steps, to accurately approximating the nonlinear integral term involving $h$. More theoretical analysis can be found in Appendix A.1 to A.3.
@@ 4.2 Datasets, Counting Hallucinations and Evaluation Protocol
The CountHalluset suite consists of three datasets for studying counting hallucinations: ToyShape, SimObject, and RealHand. Representative examples from each dataset are shown on the left side of Fig. 1. Additional statistical summaries are provided in Appendix A.4.
@@ 4.2.1 ToyShape
Counting hallucinations. According the counting criteria defined above, any image generated by the model trained on ToyShape that contains two or more shapes of the same category is considered a counting hallucination, such as an image containing two pentagons in the top-right region of Fig. 1. Besides, empty images (i.e., containing only the black background without any shapes) are also classified as counting hallucinations.
Counting model used in ToyShape. In our ToyShape generation experiments, we rely solely on a counting model to directly identify counting errors within images, without using a counting-ready indicator, as the dataset’s simplicity makes severely degraded images rare. We construct a large-scale ToyShape dataset of over 400,000 samples, with each category appearing 0–3 times per sample (i.e., $S_{c}=\{0,1,2,3\}$), to fine-tune a ResNet-50 (He et al., 2016). This setup captures a wide range of plausible scenarios, including normal samples and counting hallucinations, enabling accurate quantification. To further enhance the robustness of the counting model, we apply Gaussian noise during training. The model can achieve over 99.9% counting accuracy on generated data.
@@ 4.2.2 SimObject
Dataset description. The SimObject dataset includes 30,000 rendered images of everyday objects. It comprises three object categories: mug, apple, clock, each with 10 distinct intra-class variations. All objects are randomly placed on a wooden tabletop under fixed lighting conditions, with minimal mutual occlusion if any, as demonstrated in the middle of the left side of Fig. 1. The dataset is synthetically generated in Unreal Engine 5 (Epic Games, 2024), featuring photorealistic rendering with accurate simulation of object geometry, material attributes, lighting, and reflections. The counting criteria for this dataset are similar to those of the ToyShape dataset (see Sec. 4.2.1). Formally, we have $C=\{\emph{mug},\emph{apple},\emph{clock}\}$. For each class $c\in C$, we set $N_{c}({\bm{x}})\in S_{c}=\{0,1\}$, subject to the constraint $\sum_{c\in C}N_{c}({\bm{x}})\geq 1$.
@@ 4.2.3 RealHand
Dataset description. To investigate counting hallucinations in real-world scenarios, we construct the RealHand dataset comprising 5,050 human hand images sourced from the 11k Hands dataset (Afifi, 2019), Kaggle [^1] https://www.kaggle.com/, and Roboflow [^2] https://universe.roboflow.com/. Specifically, we aggregate raw images from these sources and remove those with blur or artifacts. Each image depicts a human hand with five fingers, capturing both dorsal (back) and palmar (front) views under varying lighting conditions and backgrounds (e.g., white surfaces, blackboards, walls), as shown in the lower-left of Fig. 1. The counting criteria specify that each image contains exactly five fingers, i.e., $C=\{\emph{finger}\}$ with $N_{c}({\bm{x}})\in S_{c}={5}$ for each $c\in C$, and $\sum_{c\in C}N_{c}({\bm{x}})=5$.
Counting hallucinations and non-counting failures. Generating realistic hand images is inherently prone to unexpected failures beyond counting hallucinations (e.g., extra or missing fingers), such as blurriness, severely deformed fingers, and visual artifacts. Therefore, we use a counting-ready indicator (i.e., a binary classifier) to identify these cases.
Counting-ready indicator and counting model used in RealHand. To separate counting hallucinations from previously mentioned failure modes, we finetune a MaxViT (Tu et al., 2022) model on 2.5k annotated generated images to serve as a counting-ready indicator aligned with human judgments. For counting, we finetune a YOLO-12 (Tian et al., 2025) model on 2k fingertip-annotated generated images. Based on human evaluation of 500 randomly sampled images by three annotators, the indicator and counting model achieve over 96% and 99% accuracy, respectively.
Quantification of counting hallucinations. Unlike the ToyShape and SimObject datasets, where all generated images are directly evaluated due to their simplicity and the rarity of severe failure cases (e.g., distorted or unrecognizable objects), the RealHand dataset requires an additional filtering step. Specifically, we first apply the counting-ready indicator to select hand images with sufficient visual quality, and only then use the counting model to predict the number of fingers. Figure 2 illustrates this distinction in detail.

### v2

@@ Abstract
Diffusion probabilistic models (DPMs) have demonstrated remarkable progress in generative tasks, such as image and video synthesis. However, they still often produce hallucinated samples (hallucinations) that conflict with real-world knowledge, such as generating an implausible duplicate cup floating beside another cup. Despite their prevalence, the lack of feasible methodologies for systematically quantifying such hallucinations hinders progress in addressing this challenge and obscures potential pathways for designing next-generation generative models under factual constraints. In this work, we bridge this gap by focusing on a specific form of hallucination, which we term counting hallucination, referring to the generation of an incorrect number of instances or structured objects, such as a hand image with six fingers, despite such patterns being absent from the training data. To this end, we construct a dataset suite CountHalluSet, with well-defined counting criteria, comprising ToyShape, SimObject, and RealHand. Using these datasets, we develop a standardized evaluation protocol for quantifying counting hallucinations, and systematically examine how different sampling conditions in DPMs, including solver type, ODE solver order, sampling steps, and initial noise, affect counting hallucination levels. Furthermore, we analyze their correlation with common evaluation metrics, such as Fréchet Inception Distance (FID), revealing widely-used perceptual metrics fail to capture counting hallucinations consistently.
@@ 1 Introduction
Diffusion probabilistic models (DPMs) (ddpm; adm; ldm), commonly referred to as diffusion model, have achieved considerable success in generative modeling, producing high-quality samples across various tasks such as text-to-image generation (cfg; ldm), audio synthesis (diffwave), video generation (videogen), 3D object generation (3dpoint; lion3dshape), and protein structure prediction (alphafold2; alphafold3). Despite these impressive advancements, diffusion models still struggle to faithfully capturing factual properties from training data, often generating outputs that violate consistency with real-world knowledge. A notable example is the generation of objects with an incorrect number of spatially intricate components, such as extra human fingers (handiffuser) or additional animal limbs (borji2023qualitative).
Recent studies have increasingly focused on identifying and analyzing common failure modes in text-guided diffusion models. For instance, borji2023qualitative systematically concluded several qualitative shortcomings in text-to-image models. Additionally, liu2024 proposed an adversarial search method to uncover undesirable behaviors and failure cases in text-to-image generation, while wu2023 employed an iterative approach to discover and mitigate spurious correlations.
More recently, hallucination, a specific failure mode in diffusion models, has garnered attention in the research community (aithal2024). Unlike previously discussed text‑prompt‑based failure modes, which can be evaluated by measuring the alignment between the input conditions and outputs (e.g., CLIP score (clipscore) and VQA score (huang2025t2i; zarei2024)), or more obvious failure modes such as severe distortions, blur, and artifacts, hallucinated samples (hallucinations) often appear visually plausible, and emerge independently of the conditioning signal, manifesting as subtle violations of factual consistency with respect to the training data, such as incorrect object counts, geometries that appear plausible but are physically impossible, or other breaches of physical laws. These fine-grained distinctions between valid samples and hallucinations present a critical challenge for current AI systems, which often fail to detect them reliably. For example, even state-of-the-art large vision-language models like GPT-4o still struggle with basic tasks like accurately counting the number of fingers in a clear hand image. To this end, establishing a standardized and well-defined protocol for reliably quantifying hallucinated samples is essential, as it provides a foundation for outlining pathways toward designing next-generation generative models with strict factual consistency, an ability that is critical for their role as world models (ha2018world; ding2024understanding; bruce2024genie).
@@
In our quantification protocol for counting hallucinations, a critical step is the accurate identification and quantification of objects. To this end, we employ counting models tailored to each dataset. Leveraging these datasets and counting models, we examine how various denoising conditions (e.g., sampling steps, initial noise, the order of ODE solvers, and solver type) within diffusion models affect counting hallucination rates. Besides, we conduct correlation analysis between counting hallucination rate and perceptual metrics, including Fréchet Inception Distance (FID) fid and MUSIQ ke2021musiq.
@@
  • Commonly applied numerical strategies in diffusion models, such as increasing sampling steps from 25 to 100, often fail to mitigate counting hallucinations in generating human hands, even though these techniques are effective in simpler synthetic datasets.
  • A weak correlation is observed between the counting hallucination rate and perceptual metrics, whereas a strong correlation emerges between the non-counting failure rate and these metrics.
  • Based on one of our findings, we propose a novel and promising training paradigm, termed joint-diffusion models, which can significantly reduce counting-based hallucinations and non-counting failures in the RealHand dataset.
@@ 2.1 Hallucinations in Diffusion Models
Inspired by the growing attention to hallucination in large language models (LLMs) (kadavath2022; ji2023survey; huang2023survey), the study of hallucinations in unconditional diffusion models has also garnered increasing interest. In particular, aithal2024 proposed a qualitative hypothesis suggesting that hallucinated samples may arise from data interpolating between distinct modes. Despite this insight into potential causes, a systematic quantitative analysis of hallucinations in diffusion models remains lacking.
@@ 2.2 Quantitative Metrics in Image Generation
The Visual Question Answering (VQA) score (huang2025t2i; zarei2024) and the CLIP score (clipscore) are two widely used metrics for evaluating the performance of text‑guided generated images. However, such evaluation methods that rely on language‑based models are inherently limited when applied to assess hallucinated images. One line of evidence is the statistical lower bound on the propensity of pretrained language models to hallucinate certain fact types, which constrains the upper bound of detection accuracy based solely on language models (llm-must-hallu). Another indication is the persistent performance gap between human annotators and LLMs in classifying hallucinated textual content (truthfulqa; halueval).
In addition, several widely used quantitative evaluation metrics, such as FID (fid), IS (is), and Precision & Recall (precision_recall), and MUSIQ (ke2021musiq) are commonly employed in image generation tasks. These metrics rely on the feature embeddings extracted from a pretrained Inception-v3 model (inception_v3). However, whether these metrics can capture the severity of hallucinations in generated datasets remains an open question. In this work, we aim to address this gap from the perspective of counting hallucinations.
@@ 3.1 Forward (Diffusion) Process
Forward Process. Given the observed data ${\bm{x}}_{0}\sim q_{0}({\bm{x}}_{0})$, Diffusion Probabilistic Models (DPMs) (sohl2015deep; ddpm) introduce a forward process (a.k.a, diffusion process) that gradually perturbs the data by adding Gaussian noise over $T$ discrete timesteps. At at each timestep $t$, the marginal distribution of the noisy sample ${\bm{x}}_{t}$ conditioned on the original data ${\bm{x}}_{0}$ follows:
@@
where $t\in\{0,1,\ldots,T\}$ and $\{\beta_{i}\}_{i=1}^{T}$ is a pre-defined noise schedule dependent on $t$, commonly chosen to be linear (ddpm) or cosine (improved_ddpm) schedules. The factor $\bar{\alpha}_{t}$ controls the signal-to-noise ratio at step $t$: $\bar{\alpha}_{0}=1$ (no noise) and $\bar{\alpha}_{T}\approx 0$ (nearly pure noise). An equivalent formulation of Eq. (1) is:
@@
As $T$ becomes sufficiently large, the marginal distribution $q_{0:T}({\bm{x}}_{T}|{\bm{x}}_{0})$ approaches a standard Gaussian distribution ${\mathcal{N}}(\bm{0},{\bm{I}})$. In widely-used DPMs (ddpm; improved_ddpm), $T$ is typically set to 1000.
@@
Training Objective. The reverse process (a.k.a., denoising process) of DPMs aims to iteratively recover the observed data based on the diffused data ${\bm{x}}_{T}$. Specifically, DPMs attempt to predict the noise $\bm{\epsilon}$} from the data ${\bm{x}}_{t}$ by using a neural network ${\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)$, known as noise prediction model. The parameter $\theta$ is optimized by minimizing the following objective (ddpm; sde):
@@
where $\nabla_{{\bm{x}}_{t}}\,log\,q_{t}({\bm{x}}_{t})$ is the only unknown term, referred as score function (sde).
@@
ODE-Based Accelerated Sampling. The ancestral sampling in Eq. (4) contains a stochastic term (kloeden1992stochastic), while there is an associated deterministic process, which can be formulated as an ordinary differential equation (ODE) (a.k.a, probability flow ODE (sde)):
@@
where ${\mathcal{G}}(t_{2},t_{1})=e^{\int_{t_{1}}^{t_{2}}f(v)\text{d}v}$ is the state transition operator for the linear part $\frac{\text{d}{\bm{z}}}{\text{d}\tau}=f(\tau){\bm{z}}$. $\Delta t>0$ denotes the step size, allowing ODE-based solvers to accelerate the sampling process with larger step sizes compared to ancestral sampling. The total number of sampling steps is given by $T/\Delta t$. Since the linear part ${\mathcal{G}}(t_{2},t_{1})$ can be computed in closed form, most accelerated ODE-based samplers (ddim; pndm; deis; dpm-solver; dpm-solver++) devote their numerical efforts, such as higher solver orders and more sampling steps, to accurately approximating the nonlinear integral term involving $h$. More theoretical analysis can be found in Appendix A.1 to A.3.
@@ 4.2 Datasets, Counting Hallucinations and Evaluation Protocol
The CountHalluset suite consists of three datasets for studying counting hallucinations: ToyShape, SimObject, and RealHand. Examples from each dataset and corresponding counting hallucinations are shown in Fig. A1 and A2. Additional statistical summaries are provided in Appendix A.4.
@@ 4.2.1 ToyShape
Counting hallucinations. According the counting criteria defined above, any image generated by the model trained on ToyShape that contains two or more shapes of the same category is considered a counting hallucination, such as an image containing two pentagons in the top-right region of Fig. 1. Besides, empty images (i.e., containing only the black background without any shapes) are also classified as counting hallucinations. Examples can be found in Appendix A.5.
Counting model used in ToyShape. We utilize a counting model to identify counting errors within images, without using a counting-ready indicator, as the dataset’s simplicity makes severely degraded images rare. In specific, we fine-tune a ResNet-50 (resnet) using over 400k ToyShape data, with each category appearing 0–3 times per sample (i.e., $S_{c}=\{0,1,2,3\}$), aiming to capture a wide range of plausible scenarios, including counting-valid samples and counting hallucinations, enabling accurate quantification. The model can achieve over 99.9% counting accuracy on generated data.
@@ 4.2.2 SimObject
Dataset description. The SimObject dataset includes 30,000 rendered images of everyday objects. It comprises three object categories: mug, apple, clock, each with 10 distinct intra-class variations. All objects are randomly placed on a wooden tabletop under fixed lighting conditions, with minimal mutual occlusion if any, as demonstrated in the middle of the left side of Fig. 1. The dataset is synthetically generated in Unreal Engine 5 (unrealengine), featuring photorealistic rendering with accurate simulation of object geometry, material attributes, lighting, and reflections. The counting criteria for this dataset are similar to those of the ToyShape dataset (see Sec. 4.2.1). Formally, we have $C=\{\emph{mug},\emph{apple},\emph{clock}\}$. For each class $c\in C$, we set $N_{c}({\bm{x}})\in S_{c}=\{0,1\}$, subject to the constraint $\sum_{c\in C}N_{c}({\bm{x}})\geq 1$.
@@ 4.2.3 RealHand
Dataset description. To investigate counting hallucinations in real-world scenarios, we construct the RealHand dataset comprising 5,050 human hand images sourced from the 11k Hands dataset (hand11k), Kaggle [^1] https://www.kaggle.com/, and Roboflow [^2] https://universe.roboflow.com/. Specifically, we aggregate raw images from these sources and remove those with blur or artifacts. Each image depicts a human hand with five fingers, capturing both dorsal (back) and palmar (front) views under varying lighting conditions and backgrounds (e.g., white surfaces, blackboards, walls), as shown in the lower-left of Fig. 1. The counting criteria specify that each image contains exactly five fingers, i.e., $C=\{\emph{finger}\}$ with $N_{c}({\bm{x}})\in S_{c}={5}$ for each $c\in C$, and $\sum_{c\in C}N_{c}({\bm{x}})=5$.
Counting hallucinations and non-counting failures. Generating realistic hand images is inherently prone to unexpected failures beyond counting hallucinations (e.g., extra or missing fingers), such as blurriness, severely deformed fingers, and visual artifacts. Therefore, we use a counting-ready indicator (i.e., a binary classifier) to identify these cases. Examples of counting hallucinations and non-counting failure samples can be found in Appendix A.7.
Counting-ready indicator and counting model used in RealHand. To separate counting hallucinations from previously mentioned failure modes, we finetune a MaxViT (tu2022maxvit) model on 2.5k annotated generated images to serve as a counting-ready indicator aligned with human judgments. For counting, we finetune a YOLO-12 (tian2025yolov12) model on 2k fingertip-annotated generated images. Based on human evaluation of 500 randomly sampled images by three annotators, the indicator and counting model achieve over 96% and 99% accuracy, respectively. Examples of YOLO detection results can be found in Appendix A.6.
Quantification of counting hallucinations. Unlike the ToyShape and SimObject datasets, where all generated images are directly evaluated due to their simplicity and the rarity of severe failure cases (e.g., distorted or unrecognizable objects), the RealHand dataset requires an additional filtering step. Specifically, we first apply the counting-ready indicator to select hand images with sufficient visual quality, and only then use the counting model to predict the number of fingers, as shown in Figure 2.